Evaluate:
xy^2(xy-x^2) If x = -1/3 and y = 1/6

Answer by checkley77(12844)

xy^2(xy-x^2) If x = -1/3 and y = 1/6
-1/3*1/6^2(-1/3*1/6-(-1/3^2))
-1/3*1/36(-1/18+1/9)
-1/108(-1+2)/18
-1/108*1/18
-1/1,944 OR .0005144 ANSWER.